OpenStudy (butterflyhope):

Order from least to greatest by using <: -3^2; 3^2; -3^-2; 3^-2 Thanks :)

OpenStudy (gorv):

negative less than positive

OpenStudy (gorv):

-3^2 and -3^-2 are minimum and 3^2 and 3^-2 will be max

OpenStudy (gorv):

now we have to arrange in order

OpenStudy (gorv):

if power is negative than take reciprocal or revese the number like x^-2=1/x^2

OpenStudy (gorv):

can u do it now ??

OpenStudy (triciaal):

one approach is to find the value of each and place on the number line (-3)^2 = (-3)(-3) = (3)^2 =(3)(3) = (-3)^-2 = (-3)^0(-3)^-2 =1/(-3)^2=1/(-3)(-3)= (3)^-2 = 1/(3)(3) =

OpenStudy (triciaal):

was that all of the question?

OpenStudy (mathstudent55):

@ButterflyHope Did you get an answer to this question?

OpenStudy (butterflyhope):

Yes sorry, thanks @triciaal and @gorv

OpenStudy (mathstudent55):

Here are the numbers in the order you listed them: \(-3^2 = -(3^2) = - 9\) \(3^2 = 9\) \(-3^{-2} = -(3^{-2}) = - \dfrac{1}{3^2} = - \dfrac{1}{9} \) \(3^{-2} = \dfrac{1}{3^2} = \dfrac{1}{9}\) The order is: \(-3^2 \lt -3^{-2} \lt 3^{-2} \lt 3^2\)